# Probabilistic enumeration and equivalence of nonisomorphic trees

### v2

@@ Abstract
Pólya trees, Otter trees
@@ 1 Introduction
$u_{n}=n^{n-2}.$ (1.1)
Numerous proofs of this quantity are known [6, 20, 18, 24, 32, 11, 1]. A common idea is to mark a root vertex. This way, each tree with $n$ vertices corresponds to precisely $n$ rooted versions, and the exponential generating series $T(z)=\sum_{n\geq 1}\frac{nu_{n}}{n!}z^{n}$ of rooted trees satisfies
@@
In order to tackle such challenges, Pólya [25] developed a general theory for counting objects up to isomorphism. In particular, he reformulated Cayley’s functional equation for the ordinary generating series $A(z)=\sum_{n\geq 1}a_{n}z^{n}$ to
@@
Later, Otter [21] derived the asymptotic formula
$a_{n}\sim c_{A}n^{-3/2}\rho^{-n}$ (1.2)
@@
A full asymptotic expansion was derived by Genitrini [12] using methods of analytic combinatorics [10, 9]. Otter additionally obtained the asymptotic number of unlabelled trees. Our first contribution in the present work is to provide a new probabilistic proof of his famous result:
Theorem 1.1 (Otter [21]).
@@
from which the asymptotic expansion of $f_{n}$ may readily be deduced. Alternative proofs of the dissymmetry equation were given by Harary [16] by using directly equations by Pólya [25], and by Bodirsky, Fusy, Kang and Vigerske [5] by introducing the cycle pointing method. Another proof of Equation (1.3) was given in [8, Thm. 13] using analytic integration and singularity analysis. The proof of Theorem 1.1 we present here is based on probabilistic methods and does not recover the dissymmetry equation.
The development of methods for asymptotically counting objects up to symmetry is of general interest since there are clear problems in enumerative combinatorics that remain open and pose a serious challenge. A prominent example is the asymptotic enumeration of unlabelled planar graphs, that is, graphs that may be drawn in the plane so that edges only intersect at endpoints. Liskovet and Walsh [19] formulated a program for their enumeration. Bender, Gao and Wormald [2] established the asymptotic number of labelled $2$-connected planar graphs, and the asymptotic number of connected and unrestricted labelled planar graphs was then obtained in the breakthrough result by Giménez and Noy [13]. An alternative combinatorial approach was later given by Chapuy, Fusy, Kang and Shoilekova [7], and a third probabilistic proof via large deviation methods was given by S. [31]. Determining the asymptotic number of nonisomorphic planar graphs with a given number of vertices however remains a major challenge that requires the development of new methods and approaches.
@@
$\mathbb{P}(F(\mathsf{A}_{n})\in\mathcal{E})\frac{a_{n}}{nf_{n}}\leq\mathbb{P}(\mathsf{F}_{n}\in\mathcal{E})\leq\mathbb{P}(F(\mathsf{A}_{n})\in\mathcal{E})\frac{a_{n}}{f_{n}}.$ (1.4)
@@
A minor drawback to this generating function approach is that we need to perform this transfer individually for each functional under consideration. A more severe drawback is that sometimes this transfer is not feasible, for example when we want to prove functional convergence of some contour function to a limiting stochastic process, or Gromov–Hausdorff–Prokhorov convergence to a limiting random real tree. This problem was noted for example by Haas and Miermont [15], who established Aldous’ Brownian tree as scaling limit of random Pólya trees, and noted that at that time the scaling limit of unlabelled unrooted trees was still an open problem. In order to address this issue, an approach that allows a transfer of “practically all” properties of $\mathsf{A}_{n}$ to $\mathsf{F}_{n}$ was presented by S. [27, Thm. 1.3]. Using the mentioned method of cycle pointing [5] it was shown that there exists a random tree $\mathsf{B}_{n}$ with stochastically bounded size $K_{n}=O_{p}(1)$ so that $\mathsf{B}_{n}$ is independent from $(\mathsf{A}_{k})_{k\geq 1}$ and such that the unrooted tree $\mathsf{A}_{n-K_{n}}+\mathsf{B}_{n}$ obtained by identifying the root vertices of $\mathsf{A}_{n-K_{n}}$ and $\mathsf{B}_{n}$ satisfies
@@ 2 Proof of main theorems
for each free tree $F\in\mathcal{F}_{n}$. This well-known fact [18] lies at the heart of the theory of cycle index series, and ensures that the exponential generating series
@@
the exponential generating series of symmetries of rooted trees that only fix the root vertex. The equation for $H(z)$ follows from the fact that any such symmetry induces a fixed point free permutation $\nu$ of the branches attached to the root, which may be grouped according to the length of the corresponding cycle of $\nu$. For any $i\geq 2$, $A(z^{i})/i$ enumerates $i$-tuples of (isomorphic) rooted trees equipped with an automorphism that permutes them cyclically, see for example [18, Prop. 10]. Hence $\exp(\sum_{i\geq 2}A(z^{i})/i)$ counts unordered collections of such objects for all $i\geq 2$, yielding the stated equation for $H(z)$.
@@
$[z^{n}]\left(F(z)-\mathrm{Sym}_{0}(\mathcal{U})(z)\right)=\frac{1}{2}\mathbb{P}(S_{N}=n)\rho^{-n}.$ (2.2)
By [4, Thm. 1, (ii), (iii)] and [17, Thm. 1.1] it follows that
@@
$[z^{n}]\left(F(z)-\mathrm{Sym}_{0}(\mathcal{U})(z)\right)\sim\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{3/2}n^{-5/2}\rho^{-n}.$ (2.3)
@@
$[z^{n}]\mathrm{Sym}_{0}(\mathcal{U})(z)\leq[z^{n}]A(z^{2})$ (2.4)
for all $n\geq 1$. Using $\rho<1$ it immediately follows from (2.3) that
$[z^{n}]F(z)\sim\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{3/2}n^{-5/2}\rho^{-n}.$ (2.5)
@@
By [4, Thm. 1, (ii), (iii)] and [17, Thm. 1.1] we have again
@@
$[z^{n}]A(z)\sim\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{1/2}n^{-3/2}\rho^{-n}.$ (2.6)
Hence, we recover Equations (1.2) and (1.3) with
$c_{A}=\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{1/2}\qquad\text{and}\qquad c_{F}=\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{3/2}.$ (2.7)
@@
The random free tree $\mathsf{F}_{n}$ may be generated by taking a uniformly selected symmetry from $\mathrm{Sym}(\mathcal{U})[\{1,\ldots,n\}]$ instead. By Equation (2.4) the probability for this symmetry to have no fixed points is given by
@@
Hence, using Equation (2.2), it follows that
@@
$\sum_{\begin{subarray}{c}0\leq k\leq n\\ |k-n/\mathbb{E}[X]|\geq n^{\alpha}\end{subarray}}\frac{[z^{n}]\mathrm{Sym}_{k}(\mathcal{U})(z)}{f_{n}}\leq\exp(-\Theta(n^{2\alpha-1})).$ (2.8)
By analogous arguments (or by combining the rough bounds (1.4) with Inequality (2.8)), a similar bound holds for the symmetries of rooted trees:
@@
For Pólya trees, bounds and limits for the number of fixed points have also been carried out in previous works [22, 26, 14]. They may also be expressed as component counts of a Gibbs partition in the dense regime [30].
@@
Using the second order asymptotics [12] for Pólya trees we obtain
$a_{n}/f_{n}=n/\mathbb{E}[X]+O(1),$ (2.9)
@@ 3.1 Trees with degree restrictions
S. [27, Thm. 1.3] proved that indeed $\mathsf{F}_{n}^{\Omega}$ admits an approximation in total variation by an $(n-O_{p}(1))$-vertex random Pólya tree with outdegrees in $\Omega^{*}$ and a small tree attached to its root to reach total vertex size $n$.
However, our assumption $\Omega\neq\mathbb{N}$ ensures that there exists an integer $k_{0}\geq 1$ with $k_{0}\in\Omega^{*}\setminus\Omega$. The limiting probability for the root of $\mathsf{A}_{n}^{\Omega^{*}}$ to have degree $k_{0}$ is positive, whereas $\mathsf{F}_{n}^{\Omega}$ by construction has no vertex with degree $k_{0}$. Thus, $\mathsf{F}_{n}^{\Omega}$ and $\mathsf{A}_{n}^{\Omega^{*}}$ are not asymptotically equivalent, and S. [27, Thm. 1.3] already provides the best comparison.
@@
Due to the strong similarity with the unrestricted case we only sketch the necessary steps. The probability that a random $\mathcal{F}^{\Omega}$-symmetry with $n$ vertices has no fixed points is exponentially small by analogous arguments as in the unrestricted case. The medium deviation concentration inequality for the number of fixed points in an ${\mathcal{A}}^{\Omega^{*}}$-symmetry was already noted in [22] and [26]. The branching process methods there may be easily adapted to get the same bound for ${\tilde{\mathcal{A}}}^{\Omega}$-symmetries. We use branching process methods here in order to elegantly avoid having to deal with the restrictions imposed by vertex degrees in the fixed point subtree to the sum of the root degrees in the attached branches. The same concentration inequality for unrooted trees then follows by using rough bounds as in (1.4). Hence the rest of the proof is then fully analogous to the unrestricted case. The required bound on the second order term in the asymptotic expansion for the number of rooted trees required to generalize Equation (2.9) follows readily by transfer theorems from the known square-root singular expansion. ∎
The branches attached to the root of $\tilde{\mathsf{A}}_{n}^{\Omega}$ form a multiset of degree restricted Pólya trees, with the total number constrained to lie in $\Omega$. The distribution of the component sizes is hence a special case of an unlabelled Gibbs partition (with all weights set to $1$). Using the main result of [28] (slightly adapted to take into account periodicities) it follows that precisely one of the branches attached to the root of $\tilde{\mathsf{A}}_{n}^{\Omega}$ is asymptotically macroscopic, with the rest having a stochastically bounded total size. The distribution of the largest branch is hence that of $\mathsf{A}_{n-K_{n}}^{\Omega^{*}}$ for some independent stochastically bounded non-negative integer $K_{n}$. In other words, $\tilde{\mathsf{A}}_{n}^{\Omega}$ (and by Proposition 3.1 also $\mathsf{T}_{n}^{\Omega}$) therefore looks like an $(n-O_{p}(1))$-vertex random Pólya tree with outdegrees in $\Omega^{*}$ and a small tree attached to its root to reach total vertex size $n$. Thus, we recover the main result of S. [27, Thm. 1.3] in a different way without using the cycle pointing method.
@@ 3.2 Subcritical classes of graphs
The definition of a subcritical class of unlabelled graphs [8, Def. 10] involves the concept of cycle index series $Z_{G}$ of a graph class $\mathcal{G}$, as well as operations such as rooted and derived classes of structures. The cycle index series of $\mathcal{G}$ is a power series in countably infinitely many variables
@@
with symmetries of graphs being defined analogously as for trees, and $\sigma_{i}$ denoting the number of cycles of length $i$ in $\sigma$ for all $i\geq 1$. Likewise, for rooted graphs all symmetries are required to fix the root vertex, and the cycle index series for classes of rooted graphs are defined accordingly. Derived graphs are rooted graphs, except that the root-vertex is required to have a place-holder label *. That is, for a derived class of graphs the symmetries on a set of vertices $U$ are pairs $(G,\sigma)$ with $G$ a rooted graph on the vertex set $U\cup\{*\}$ (assuming always $\notin U$), and $\sigma:U\to U$ a permutation of $U$ such that the canonical bijective extension $U\cup\{*\}\to U\cup\{*\}$ is a graph automorphism of $G$. We refer the reader to [18, 3] for details on the technical background of these notions.
@@
The class is termed subcritical in the unlabelled case [8, Def. 10], if additionally the following conditions are met.
@@
Subcritical classes of graphs are sometimes called tree-like, as they admit the Brownian tree as scaling limit and therefore their global shape resembles that of a tree. This was shown for the labelled case in [23], the unlabelled rooted case in [26], and the unlabelled unrooted case in [29].
@@
An earlier approximation result [29, Thm. 1] formulated for block-weighted showed that $\mathsf{C}_{n}$ may be approximated in total variation by an $(n-O_{p}(1))$-sized random rooted graph with a small graph attached to its root to make up for the missing number of vertices. While this is already sufficient for all practical purposes, the following approximation result provides a more elegant comparison:
@@
Consequently, the fourth condition in the definition of subcriticality ensures that only an exponentially small fraction of symmetries of $\mathcal{C}$ have no fixed-points. Furthermore, the concentration inequality for the number of fixed points in a symmetry of $\mathcal{C}^{\bullet}$ was already obtained in [26] via branching process methods. The bound on the second order term in the asymptotics for the number of rooted graphs required to generalize Equation (2.9) follows readily from the square-root singular expansions given in [8]. Thus, the proof may be carried out entirely analogously as the proof of Theorem 1.2. ∎
Acknowledgement
I warmly thank the referees for their helpful remarks.

### v3

@@ Abstract
Unlabelled trees, Asymptotic Enumeration
@@ 1 Introduction
$u_{n}=n^{n-2}.$ (1)
Numerous proofs of this quantity are known Borchardt (1860); Moon (1967); Joyal (1981); Pitman (1999); van der Hofstad (2017); Foata and Fuchs (1970); Addario-Berry et al. (2023). A common idea is to mark a root vertex. This way, each tree with $n$ vertices corresponds to precisely $n$ rooted versions, and the exponential generating series $T(z)=\sum_{n\geq 1}\frac{nu_{n}}{n!}z^{n}$ of rooted trees satisfies
@@
In order to tackle such challenges, Pólya Pólya (1937) developed a general theory for counting objects up to isomorphism. In particular, he reformulated Cayley’s functional equation for the ordinary generating series $A(z)=\sum_{n\geq 1}a_{n}z^{n}$ to
@@
Later, Otter Otter (1948) derived the asymptotic formula
$a_{n}\sim c_{A}n^{-3/2}\rho^{-n}$ (2)
@@
A full asymptotic expansion was derived by Genitrini Genitrini (2016) using methods of analytic combinatorics Flajolet and Sedgewick (2009); Flajolet and Odlyzko (1990). Otter additionally obtained the asymptotic number of unlabelled trees. Our first contribution in the present work is to provide a new probabilistic proof of his famous result:
Theorem 1.1 (Otter Otter (1948)).
@@
from which the asymptotic expansion of $f_{n}$ may readily be deduced. Alternative proofs of the dissymmetry equation were given by Harary Harary (1955/56) by using directly equations by Pólya Pólya (1937), and by Bodirsky, Fusy, Kang and Vigerske Bodirsky et al. (2011) by introducing the cycle pointing method. Another proof of Equation (3) was given in (Drmota et al., 2011, Thm. 13) using analytic integration and singularity analysis. The proof of Theorem 1.1 we present here is based on probabilistic methods and does not recover the dissymmetry equation.
The development of methods for asymptotically counting objects up to symmetry is of general interest since there are clear problems in enumerative combinatorics that remain open and pose a serious challenge. A prominent example is the asymptotic enumeration of unlabelled planar graphs, that is, graphs that may be drawn in the plane so that edges only intersect at endpoints. Liskovet and Walsh Liskovets and Walsh (1987) formulated a program for their enumeration. Bender, Gao and Wormald Bender et al. (2002) established the asymptotic number of labelled $2$-connected planar graphs, and the asymptotic number of connected and unrestricted labelled planar graphs was then obtained in the breakthrough result by Giménez and Noy Giménez and Noy (2009). An alternative combinatorial approach was later given by Chapuy, Fusy, Kang and Shoilekova Chapuy et al. (2008), and a third probabilistic proof via large deviation methods was given by S. Stufler (2023). Determining the asymptotic number of nonisomorphic planar graphs with a given number of vertices however remains a major challenge that requires the development of new methods and approaches.
@@
$\mathbb{P}(F(\mathsf{A}_{n})\in\mathcal{E})\frac{a_{n}}{nf_{n}}\leq\mathbb{P}(\mathsf{F}_{n}\in\mathcal{E})\leq\mathbb{P}(F(\mathsf{A}_{n})\in\mathcal{E})\frac{a_{n}}{f_{n}}.$ (4)
@@
A minor drawback to this generating function approach is that we need to perform this transfer individually for each functional under consideration. A more severe drawback is that sometimes this transfer is not feasible, for example when we want to prove functional convergence of some contour function to a limiting stochastic process, or Gromov–Hausdorff–Prokhorov convergence to a limiting random real tree. This problem was noted for example by Haas and Miermont Haas and Miermont (2012), who established Aldous’ Brownian tree as scaling limit of random Pólya trees, and noted that at that time the scaling limit of unlabelled unrooted trees was still an open problem. In order to address this issue, an approach that allows a transfer of “practically all” properties of $\mathsf{A}_{n}$ to $\mathsf{F}_{n}$ was presented by S. (Stufler, 2019, Thm. 1.3). Using the mentioned method of cycle pointing Bodirsky et al. (2011) it was shown that there exists a random tree $\mathsf{B}_{n}$ with stochastically bounded size $K_{n}=O_{p}(1)$ so that $\mathsf{B}_{n}$ is independent from $(\mathsf{A}_{k})_{k\geq 1}$ and such that the unrooted tree $\mathsf{A}_{n-K_{n}}+\mathsf{B}_{n}$ obtained by identifying the root vertices of $\mathsf{A}_{n-K_{n}}$ and $\mathsf{B}_{n}$ satisfies
@@ 2 Proof of main theorems
for each free tree $F\in\mathcal{F}_{n}$. This well-known fact Joyal (1981) lies at the heart of the theory of cycle index series, and ensures that the exponential generating series
@@
the exponential generating series of symmetries of rooted trees that only fix the root vertex. The equation for $H(z)$ follows from the fact that any such symmetry induces a fixed point free permutation $\nu$ of the branches attached to the root, which may be grouped according to the length of the corresponding cycle of $\nu$. For any $i\geq 2$, $A(z^{i})/i$ enumerates $i$-tuples of (isomorphic) rooted trees equipped with an automorphism that permutes them cyclically, see for example (Joyal, 1981, Prop. 10). Hence $\exp(\sum_{i\geq 2}A(z^{i})/i)$ counts unordered collections of such objects for all $i\geq 2$, yielding the stated equation for $H(z)$.
@@
$[z^{n}]\left(F(z)-\mathrm{Sym}_{0}(\mathcal{U})(z)\right)=\frac{1}{2}\mathbb{P}(S_{N}=n)\rho^{-n}.$ (6)
By (Bloznelis, 2019, Thm. 1, (ii), (iii)) and (Hilberdink, 1996, Thm. 1.1) it follows that
@@
$[z^{n}]\left(F(z)-\mathrm{Sym}_{0}(\mathcal{U})(z)\right)\sim\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{3/2}n^{-5/2}\rho^{-n}.$ (7)
@@
$[z^{n}]\mathrm{Sym}_{0}(\mathcal{U})(z)\leq[z^{n}]A(z^{2})$ (8)
for all $n\geq 1$. Using $\rho<1$ it immediately follows from (7) that
$[z^{n}]F(z)\sim\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{3/2}n^{-5/2}\rho^{-n}.$ (9)
@@
By (Bloznelis, 2019, Thm. 1, (ii), (iii)) and (Hilberdink, 1996, Thm. 1.1) we have again
@@
$[z^{n}]A(z)\sim\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{1/2}n^{-3/2}\rho^{-n}.$ (10)
Hence, we recover Equations (2) and (3) with
$c_{A}=\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{1/2}\qquad\text{and}\qquad c_{F}=\frac{1}{\sqrt{2\pi}}\mathbb{E}[X]^{3/2}.$ (11)
@@
The random free tree $\mathsf{F}_{n}$ may be generated by taking a uniformly selected symmetry from the set $\mathrm{Sym}(\mathcal{U})[\{1,\ldots,n\}]$ instead. By Equation (8) the probability for this symmetry to have no fixed points is given by
@@
Hence, using Equation (6), it follows that
@@
$\sum_{\begin{subarray}{c}0\leq k\leq n\\ |k-n/\mathbb{E}[X]|\geq n^{\alpha}\end{subarray}}\frac{[z^{n}]\mathrm{Sym}_{k}(\mathcal{U})(z)}{f_{n}}\leq\exp(-\Theta(n^{2\alpha-1})).$ (12)
By analogous arguments (or by combining the rough bounds (4) with Inequality (12)), a similar bound holds for the symmetries of rooted trees:
@@
For Pólya trees, bounds and limits for the number of fixed points have also been carried out in previous works Panagiotou and Stufler (2018); Stufler (2018); Gittenberger et al. (2018). They may also be expressed as component counts of a Gibbs partition in the dense regime Stufler (2024).
@@
Using the second order asymptotics Genitrini (2016) for Pólya trees we obtain
$a_{n}/f_{n}=n/\mathbb{E}[X]+O(1),$ (13)
@@ 3.1 Trees with degree restrictions
S. (Stufler, 2019, Thm. 1.3) proved that indeed $\mathsf{F}_{n}^{\Omega}$ admits an approximation in total variation by an $(n-O_{p}(1))$-vertex random Pólya tree with outdegrees in $\Omega^{*}$ and a small tree attached to its root to reach total vertex size $n$.
However, our assumption $\Omega\neq\mathbb{N}$ ensures that there exists an integer $k_{0}\geq 1$ with $k_{0}\in\Omega^{*}\setminus\Omega$. The limiting probability for the root of $\mathsf{A}_{n}^{\Omega^{*}}$ to have degree $k_{0}$ is positive, whereas $\mathsf{F}_{n}^{\Omega}$ by construction has no vertex with degree $k_{0}$. Thus, $\mathsf{F}_{n}^{\Omega}$ and $\mathsf{A}_{n}^{\Omega^{*}}$ are not asymptotically equivalent, and S. (Stufler, 2019, Thm. 1.3) already provides the best comparison.
@@
Due to the strong similarity with the unrestricted case we only sketch the necessary steps. The probability that a random $\mathcal{F}^{\Omega}$-symmetry with $n$ vertices has no fixed points is exponentially small by analogous arguments as in the unrestricted case. The medium deviation concentration inequality for the number of fixed points in an ${\mathcal{A}}^{\Omega^{*}}$-symmetry was already noted in Panagiotou and Stufler (2018) and Stufler (2018). The branching process methods there may be easily adapted to get the same bound for ${\tilde{\mathcal{A}}}^{\Omega}$-symmetries. We use branching process methods here in order to elegantly avoid having to deal with the restrictions imposed by vertex degrees in the fixed point subtree to the sum of the root degrees in the attached branches. The same concentration inequality for unrooted trees then follows by using rough bounds as in (4). Hence the rest of the proof is then fully analogous to the unrestricted case. The required bound on the second order term in the asymptotic expansion for the number of rooted trees required to generalize Equation (13) follows readily by transfer theorems from the known square-root singular expansion. ∎
The branches attached to the root of $\tilde{\mathsf{A}}_{n}^{\Omega}$ form a multiset of degree restricted Pólya trees, with the total number constrained to lie in $\Omega$. The distribution of the component sizes is hence a special case of an unlabelled Gibbs partition (with all weights set to $1$). Using the main result of Stufler (2020) (slightly adapted to take into account periodicities) it follows that precisely one of the branches attached to the root of $\tilde{\mathsf{A}}_{n}^{\Omega}$ is asymptotically macroscopic, with the rest having a stochastically bounded total size. The distribution of the largest branch is hence that of $\mathsf{A}_{n-K_{n}}^{\Omega^{*}}$ for some independent stochastically bounded non-negative integer $K_{n}$. In other words, $\tilde{\mathsf{A}}_{n}^{\Omega}$ (and by Proposition 3.1 also $\mathsf{T}_{n}^{\Omega}$) therefore looks like an $(n-O_{p}(1))$-vertex random Pólya tree with outdegrees in $\Omega^{*}$ and a small tree attached to its root to reach total vertex size $n$. Thus, we recover the main result of S. (Stufler, 2019, Thm. 1.3) in a different way without using the cycle pointing method.
@@ 3.2 Subcritical classes of graphs
The definition of a subcritical class of unlabelled graphs (Drmota et al., 2011, Def. 10) involves the concept of cycle index series $Z_{G}$ of a graph class $\mathcal{G}$, as well as operations such as rooted and derived classes of structures. The cycle index series of $\mathcal{G}$ is a power series in countably infinitely many variables
@@
with symmetries of graphs being defined analogously as for trees, and $\sigma_{i}$ denoting the number of cycles of length $i$ in $\sigma$ for all $i\geq 1$. Likewise, for rooted graphs all symmetries are required to fix the root vertex, and the cycle index series for classes of rooted graphs are defined accordingly. Derived graphs are rooted graphs, except that the root-vertex is required to have a place-holder label *. That is, for a derived class of graphs the symmetries on a set of vertices $U$ are pairs $(G,\sigma)$ with $G$ a rooted graph on the vertex set $U\cup\{*\}$ (assuming always $\notin U$), and $\sigma:U\to U$ a permutation of $U$ such that the canonical bijective extension $U\cup\{*\}\to U\cup\{*\}$ is a graph automorphism of $G$. We refer the reader to Joyal (1981); Bergeron et al. (1998) for details on the technical background of these notions.
@@
The class is termed subcritical in the unlabelled case (Drmota et al., 2011, Def. 10), if additionally the following conditions are met.
@@
Subcritical classes of graphs are sometimes called tree-like, as they admit the Brownian tree as scaling limit and therefore their global shape resembles that of a tree. This was shown for the labelled case in Panagiotou et al. (2016), the unlabelled rooted case in Stufler (2018), and the unlabelled unrooted case in Stufler (2021).
@@
An earlier approximation result (Stufler, 2021, Thm. 1) formulated for block-weighted showed that $\mathsf{C}_{n}$ may be approximated in total variation by an $(n-O_{p}(1))$-sized random rooted graph with a small graph attached to its root to make up for the missing number of vertices. While this is already sufficient for all practical purposes, the following approximation result provides a more elegant comparison:
@@
Consequently, the fourth condition in the definition of subcriticality ensures that only an exponentially small fraction of symmetries of $\mathcal{C}$ have no fixed-points. Furthermore, the concentration inequality for the number of fixed points in a symmetry of $\mathcal{C}^{\bullet}$ was already obtained in Stufler (2018) via branching process methods. The bound on the second order term in the asymptotics for the number of rooted graphs required to generalize Equation (13) follows readily from the square-root singular expansions given in Drmota et al. (2011). Thus, the proof may be carried out entirely analogously as the proof of Theorem 1.2. ∎
@@
Acknowledgements.
I warmly thank the referees for their helpful remarks. This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/PAT6732623]. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.